# Discovering EV Charging Site Archetypes Through Few Shot Forecasting: The First U.S.-Wide Study

### v1

@@ 1 Introduction
The global transportation sector is undergoing a paradigm shift, with electric vehicles (EVs) central to achieving net zero emissions in the United States [1, 2]. Achieving this transition requires not only rapid deployment of charging infrastructure, but also a nuanced understanding of EV charging behavior. Such insights directly inform investment decisions [3], grid stability [4, 5], improve energy management  [6] and dynamic pricing strategies [7, 8], and shape policies [9, 10], allowing a reliable and cost-effective transition to electric mobility. Forecasting charging demand is therefore not only a matter of operational efficiency, but also a critical lever for accelerating decarbonization.
However, despite growing interest [11], three persistent challenges remain. First, most prior work relies on small-scale datasets that do not capture the intricate complexity of real-world charging patterns across locations and usage scenarios [12, 13, 14]. Second, existing models often treat charging sites in isolation [12], forecast regional demand  [15, 16, 17, 18], or cluster by proximity [19], unable to model complex temporal and spatial relationships between sites. Third, there is limited understanding of how global knowledge can be transferred to forecast demand at newly deployed sites where there is little to no operational history [20]. Inaccurate demand modeling not only inflates operational costs, but also slows EV adoption—challenges that can be mitigated through informed interventions such as curtailment, peak-shaving through dynamic pricing, and optimized battery dispatch.
@@ 2 Methodology
To achieve this, we characterize site-level demand by aggregating into weekly utilization profiles. In addition, we extract a set of canonical features (e.g., distributional characteristics, autocorrelation structure, and outlier dynamics), summarizing key statistical properties of each time series [21]. Using these representations, we apply k-means [22] to group sites into clusters of similar behavior. For each candidate number of clusters $k\in\{1,\dots,20\}$, cluster assignments are used to train a Temporal Fusion Transformer [23] model specialized for that cluster, resulting in a set of forecasting models $\Theta=\{\theta_{1},\dots,\theta_{k}\}$. The performance of the model is then evaluated across $k$ to identify the optimal clustering. Symmetric Mean Absolute Percentage Error (sMAPE) and Root Mean Squared Error (RMSE) are used to capture complementary aspects of predictive performance, where sMAPE measures relative accuracy across scales, and RMSE emphasizes magnitude and sensitivity to large deviations. Note that similar performance gains are observed with other architectures such as TCN [24] and N-BEATS [25] with varying baseline (or global) performance.
@@ A.1 Cluster spread across regions
Figure 4 visualizes the geospatial distribution of charging site archetypes across the US. Individual sites are aggregated into hexagonal cells [26], and each hexagonal is colored by the dominant utilization cluster within its boundary. The result highlights regional demand patterns, from dense urban corridors to steady retail hubs and emerging suburban clusters. By abstracting sites into cluster-level patterns, the map provides a high-level view of infrastructure heterogeneity that informs siting, forecasting, and pricing strategies.

### v2

@@ 1 Introduction
The global transportation sector is undergoing a paradigm shift, with electric vehicles (EVs) central to achieving net zero emissions in the United States [ahmadi2019environmental, woody2023decarbonization]. Achieving this transition requires not only rapid deployment of charging infrastructure, but also a nuanced understanding of EV charging behavior. Such insights directly inform investment decisions [linjuan2024site], grid stability [dharmakeerthi2014impact, calearo2019grid], improve energy management  [huang2024unveiling] and dynamic pricing strategies [kazemtarghi2024dynamic, bernard2025dynamic], and shape policies [mastoi2022depth, li2016electric], allowing a reliable and cost-effective transition to electric mobility. Forecasting charging demand is therefore not only a matter of operational efficiency, but also a critical lever for accelerating decarbonization.
However, despite growing interest [yousuf2024depth], three persistent challenges remain. First, most prior work relies on small-scale datasets that do not capture the intricate complexity of real-world charging patterns across locations and usage scenarios [van2024large, martins2022short, huttel2021deep]. Second, existing models often treat charging sites in isolation [van2024large], forecast regional demand  [louie2017time, yi2022electric, wang2018application, kim2021forecasting], or cluster by proximity [miltner2024towards], unable to model complex temporal and spatial relationships between sites. Third, there is limited understanding of how global knowledge can be transferred to forecast demand at newly deployed sites where there is little to no operational history [rashid2024comprehensive]. Inaccurate demand modeling not only inflates operational costs, but also slows EV adoption—challenges that can be mitigated through informed interventions such as curtailment, peak-shaving through dynamic pricing, and optimized battery dispatch.
@@ 2 Methodology
To achieve this, we characterize site-level demand by aggregating into weekly utilization profiles. In addition, we extract a set of canonical features (e.g., distributional characteristics, autocorrelation structure, and outlier dynamics), summarizing key statistical properties of each time series [Lubba2019:Catch22CAnonicalTimeseries]. Using these representations, we apply k-means [lloyd1982least] to group sites into clusters of similar behavior. For each candidate number of clusters $k\in\{1,\dots,20\}$, cluster assignments are used to train a Temporal Fusion Transformer [lim2021temporal] model specialized for that cluster, resulting in a set of forecasting models $\Theta=\{\theta_{1},\dots,\theta_{k}\}$. The performance of the model is then evaluated across $k$ to identify the optimal clustering. Symmetric Mean Absolute Percentage Error (sMAPE) and Root Mean Squared Error (RMSE) are used to capture complementary aspects of predictive performance, where sMAPE measures relative accuracy across scales, and RMSE emphasizes magnitude and sensitivity to large deviations. Note that similar performance gains are observed with other architectures such as TCN [bai2018empirical] and N-BEATS [oreshkin2019n] with varying baseline (or global) performance.
@@ A.1 Cluster spread across regions
Figure 4 visualizes the geospatial distribution of charging site archetypes across the US. Individual sites are aggregated into hexagonal cells [uber_h3_docs], and each hexagonal is colored by the dominant utilization cluster within its boundary. The result highlights regional demand patterns, from dense urban corridors to steady retail hubs and emerging suburban clusters. By abstracting sites into cluster-level patterns, the map provides a high-level view of infrastructure heterogeneity that informs siting, forecasting, and pricing strategies.